# Data Science Principles for Interpretable and Explainable AI

### v1

@@ 2.2.1 Sparse logistic regression
where . This optimization identifies a sparse coefficient vector that maximizes the log-likelihood of pairs . The parameter controls sparsity: higher values yield simpler but less expressive models. The combination of sparsity and linearity ensures interpretability. Sparsity allows many features within to be ignored during prediction, while linearity ensures that for features that do contribute, they enter the model transparently. Specifically, a one-unit increase in changes the relative risk for class by .
@@ 2.2.2 Decision trees
In contrast, decision trees do not solve a single optimization problem. Instead, they are defined through a recursive algorithm. Initially, we scan all dimensions of and determine thresholds such that maximizes classification accuracy. From these candidates, we select the optimal , defining the first partition along the axis . The two elements of the partition correspond to two leaves of a tree whose only split according to . Next, for a tree with leaves, we iteratively split each element along axis at a threshold , for all pairs of leaves and dimensions , selecting the split that maximizes accuracy. This process stops once model complexity outweighs performance on a holdout set.
Interpretability here stems from parsimony and simulatability. The final tree has only as many partition elements as leaves, and with sufficient cost on model complexity, this can be easily shown as a single tree. Further, the axis-aligned splits yield straightforward descriptions for each partition element, recoverable by tracing a path from the root to the leaf. Finally, predictions remain constant within each element, changing only across partition boundaries.
@@ 2.2.3 Advances
Sparse and tree-based models are the foundation for many proposals for interpretable machine learning. For example, Zeng et al. (2016) introduced fast decision tree algorithms that match black box algorithms in recidivism prediction, a domain with significant moral ramifications where interpretability is crucial. Similarly, Hazimeh et al. (2020) designed a decision tree variant that competes effectively with deep learning in various operations research tasks. By leveraging advances in combinatorial optimization, their approach allows interpretable trees to be learned from very large datasets without excessive compute demands.
Similarly, regularization is central to many modern interpretability workflows, even in problems that initially seem distant from traditional statistical learning frameworks. For example, Fridovich-Keil et al. (2022) developed an interpretable approach to the view synthesis problem, where many two-dimensional views of an object are consolidated into a coherent three-dimensional reconstruction. Instead of the typical deep encoder-decoder architecture, they used -regularized regression on the appropriate Fourier representation. This approach achieved better accuracy than competing deep learning approaches with far less sensitivity to hyperparameter selection, and could be trained at a fraction of the compute cost.
@@ 2.2.4 Application to the simulation
We next apply these directly interpretable models to the simulated trajectories from Section 2.1. We explore two approaches. First, we concatenate all time points for each participant into a single -dimensional vector. This method can uncover temporally localized predictive signals. Second, we derive time series summaries for each taxon, reducing the number of predictors to for different summaries. In our study, we focus solely on linear trends and curvature. For each subject and taxon , we fit a linear model , with time as a predictor. The estimated summarizes the linear trend for that trajectory. Similarly, we use the sum-of-squares of the temporal second differences to summarize curvature: . Note that summarization derives potentially relevant features that are not simple linear combinations of the original inputs.
@@
The holdout accuracy for the sparse logistic regression and decision tree models under these two representations of is given in Table 2. Regularization and tree complexity parameters were chosen using four fold cross validation. For these data, the sparse logistic regression model outperforms the decision tree. Moreover, manually derived summary statistics yield better performance compared to concatenating all inputs. This suggests a general takeaway: a simple model on the appropriate representation can achieve good performance, and performance is contingent on the representation. The widely invoked “interpretability-accuracy tradeoff” overlooks this nuance. Further, Table 2 includes the training time for the algorithms, illustrating a broad range of runtimes among interpretable models. Simple doesn’t always mean fast or scalable.
@@
With respect to interpretability, on concatenated inputs, the sparse logistic regression model with the optimal required 37 taxon-by-time point features, while the decision tree needed 13 splits. In contrast, with the summary statistic featurization, the optimal sparse logistic regression model used 50 features – the problem is lower dimensional but the signal is more enriched, a fact suggested by the difference in prediction performance. Surprisingly, this enriched signal is not accessible to the decision tree, which produces a tree with only two splits.
Data | Model | In-Sample Accuracy | Out-of-Sample Accuracy | Training Time (s) |
Original | Lasso | 81.2% | 78.0% | 3.9 |
Decision Tree | 91.6% | 70.2% | 95.9 | |
Transformer | 98.7% | 83.2% | 113.2 | |
Concept Bottleneck | 99.7% | 84.0% | 120.9 | |
Featurized | Lasso | 92.6% | 87.8% | 0.4 |
Decision Tree | 74.4% | 69.6% | 1.8 |
@@
The parsimony of these estimates aligns with our discussion of the characteristics of directly interpretable models. In both approaches, prediction relies only on a small subset of the original inputs. But how much can we trust these features? Fig 2(e - f) shows the estimated for nonzero features obtained from separately fitting the sparse logistic regression approaches on independent splits. When using concatenated inputs, only two features overlap, while with summary statistics, this increases to 18. Therefore, while the model on raw time points may be more parsimonious, it is less stable. In this sense, despite its parsimony, the first model does not lend itself naturally to interpretation. Conversely, the second appears more accurate and stable. Moreover, its features, like curvature and trends, match the existing vocabulary of domain experts, making it the more interpretable model in this context. The larger lesson is that interpretability refers to a constellation of model properties – parsimony and simulatability but also accuracy, stability, and contextual relevance – and it is worthwhile to think in these more refined terms rather than simply “interpretable” and “black box” models.
@@ 2.3 Deep learning and XAI
The sparse logistic regression with derived summary statistics seems to be our best option in the case study. However, we should be concerned by the fact that we manually define our features – what other potentially predictive features might we be missing out on? Black box algorithms are appealing because they automate the representation learning process (Bengio et al., 2013; Bengio, 2009). For example, in computer vision, deep learning models can identify higher-order, semantically meaningful features from raw pixel inputs (Zeiler and Fergus, 2013; Yosinski et al., 2015; Simonyan et al., 2014). The can learn features that activate on images of eyes simply by being given training data of faces and an appropriate loss function. Applying deep learning model to our microbiome data could reveal predictive temporal and taxonomic features (e.g., the presence of particular subcommunities) that we had not previously considered. The challenge, then, would be to explain these models in a way that advances scientific knowledge and ensures accurate medical diagnoses. To demonstrate the process, let’s explore using a transformer encoder model in our simulation.
@@ 2.3.1 Transformers
Specifically, consider a sequence of tokens . These tokens can be embedded into a sequence of high-dimensional (e.g., 512) real-valued vectors. In a word-based language model, could be one-hot encodings of the words from the language’s vocabulary, while in a vision problem, might represent neighboring patches from an image. The vectors are meant to capture the high-level features of the original token (word or patch) within its larger context (sentence or image). Therefore, they are much more valuable for downstream analysis than word identity or pixel values considered in isolation.
@@ 2.3.2 Embedding visualization
For example (Coenen et al., 2019) used embeddings to analyze the BERT language model (Devlin et al., 2019). In one experiment, they examined sentences containing words that can have multiple meanings depending on context, like, “fair.” By computing embeddings for instances of ”fair” within these sentences, it was observed that although the word remained the same, its embeddings distinctly clustered in a two-dimensional principal components projection. This clustering reflected the context dependence of the word’s meaning, with clusters representing the legal (“fair use”), mathematical (“fair coin”), and civic (“world’s fair”) uses. This qualitative study can also be accompanied by quantitative analysis. For example, Coenen et al. (2019) and Hewitt and Manning (2019) used linear probes to quantify the extent to which grammatically structure (represented by parse tree representations) could be reconstructed from BERT embeddings.
@@ 2.3.3 Integrated gradients
When seeking local explanations, embedding visualization lack sufficient specificity. We often have questions in mind that require more than model-based summaries – for example, what was it about a hotel review that led to its classification as negative sentiment, or how could a loan applicant modify their application so that it can be approved? Addressing these questions requires thinking carefully about the model’s sensitivity to perturbation. Many methods formalize this intuition – influence functions (Koh and Liang, 2017), shapely values (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016), GradCAM (Selvaraju et al., 2017), and Integrated Gradients (IG) (Sundararajan et al., 2017) for example. Here, we dive into the mechanics of IG.
@@
where is a reference sample (e.g., an all-black “empty” image) and is the model’s predicted probability for the input being assigned to class . The term represents how an infinitesimal change in the coordinate of affects the probability of assigning the input to class . At first, this appears to be sufficient for local explanation – large gradients correspond to words or pixels with the largest influence on the class assignment. However, this often fails in practice due to saturation. Specifically, deep learning classifiers often use the multiclass logistic function in the final layer, and for large input magnitudes, this function’s gradients approach zero, rendering explanations based on unsatisfactory.
@@
Integrated gradients and its local explanation counterparts are widely used in XAI, but they are also among the most controversial. Both theoretical and empirical results suggest that explanations do not necessarily respect expectations. For example, Adebayo et al. (2018) observed that saliency maps, including those from IG, derived from randomly initialized neural network classifiers look closely resemble those from properly trained ones. A theoretical study by Bilodeau et al. (2024) found that locally similar functions could nonetheless be assigned different explanations at their query points, indicating that global properties can influence local explanations. These alarming findings have renewed interest in the interpretability community to develop local explainability techniques with formal guarantees.
@@ 2.3.4 Concept bottleneck models
Both embedding analysis and integrated gradients rely solely on the model to be explained and the original training data. This is both a strength and a limitation. On the one hand, it ensures generality, the same explanation method can be effective across various applications. On the other, it limits the incorporation of domain-specific language into the explanation. Concept Bottleneck Models (CBMs) (Koh et al., 2020) offer an abstraction for infusing domain-specific language into models. CBM “concepts” bridge model and human representations of relevant data features. The overarching goal of this class of models is to create representations that can be manipulated by both the model (as high-dimensional, numerical vectors) and experts (as community-derived concepts).
@@
On a test sample , predictions are made using and . Note that concept labels are not needed at this stage. Typically, a simple model , like multiclass logistic regression, is used, This model essentially automates human-interpretable feature extraction, bridging manual feature engineering with black box representation learning. While manually designing a “wing color” feature extractor would be tedious, it is helpful to know that the model learns a proxy for such a feature, and its output can be used like any human engineered feature. Moreover, this approach streamlines counterfactual reasoning. For example, we can ask what the model would have predicted had the wing been yellow instead of red, simply by intervening on the associated with wing color. This is important in fairness or recourse analysis. For example, if we know that a feature should not be used for prediction (e.g., the race of a loan applicant), then we can zero out the coefficient from that concept in the model .
@@ 2.3.5 Application to the simulation
With these XAI techniques, we can apply black box models to our microbiome case study and still draw scientific insight from them. The simulated dataset is well-suited to a transformer encoder architecture, where each sample corresponds to a word and the full trajectory to a sentence. The final healthy vs. disease classification is analogous to text sentiment classification. Guided by these parallels, we trained a GPT-2 architecture with a random sample of 375 simulated participants; the remaining 125 are reserved for validation. We used default hyperparameters from a public implementation with the two modifications. First, we changed the input dimension to 250 to one-hot encode the species. Second, we reduced the number of layers to 6 from the default of 12, since the sample size is relatively small. After training this model for 70 epochs, the model achieves a holdout accuracy of 83.2%. This performance is noteworthy considering that it makes progress towards the best model with handcrafted features (cf. Table 2), which were designed with the true generative mechanism in mind. Indeed, considering the close alignment between handcrafted features and the generative mechanism, we suspect that the Bayes error for this problem is close to 13%. The ability to nearly match a handcrafted model underscores the appeal of modern deep learning. It is not necessarily about achieving better performance than interpretable, handcrafted counterparts on all problems; rather, it comes from the potential to bypasses feature engineering – a traditionally labor-intensive and time-consuming step – while achieving comparable performance.
@@
The remaining challenge lies in understanding how the model achieved this performance. We consider each of the techniques discussed above. Fig 3 offers two global views of the data. Fig 3a, gives the sparse PCA projections derived from the scaled matrix of community profiles over time, while Fig 3b shows the analog for the transformer representations. The embeddings in (b) more clearly distinguish healthy and diseased subjects, capturing more of the total variance within the top two dimensions. This is as expected, since the final classification uses a linear hyperplane, encouraging the transformer to learn a representation that ensures linearly separability. To better understand the features encoded by the transformer, we focus on the subset of embeddings associated with species 21. Fig 3c shows sparse PCA applied to solely these embeddings. We have highlighted samples that lie near the linear interpolation between samples 110 and 378 in the original embedding space, and their original trajectories are shown in panel (d). Along this interpolation, the peaks and valleys become less pronounced, suggesting that the model has learned to differentiate curves with different trajectory. This is a subtle feature, and one that we know by design is related to the response. In contrast, some visually prominent features unrelated to disease, like the maximum value of the series, are appropriately disregarded.
@@
We next study local explanations using IG. Based on the input structure, IG returns a value for each time point-by-taxon-by-taxon combination. Example values are given in Fig 4. In each panel, we can identify time points that informed the final classification and the direction in which perturbations change the target class label. “+” denotes that an increased abundance at that time point increases the probability of correct classification; “-” suggests the opposite. Larger symbols imply larger changes in probability. For example, for subject 6, the first ten time points and the valley surrounding time 20 were most important. Increases in either of these intervals increase the probability of correctness. In contrast, for subject 63, an increase in the valley surrounding time 20 decreases this probability. Since these subjects are from opposite classes, deeper valleys around time 20 seem to be related to health. However, beware that these conclusions are essentially local, and this pattern need not hold across all samples for IG to declare it important. In general, IG effectively localizes the responsibility for a classified instance down to a subset of time points, aiding error analysis and highlighting subsets of features that could flip the class of a prediction.
@@
Next, we take a closer look at the use of CBMs in our case study. Recall that concepts are a form of dense annotation for each sample. We extend our original generative mechanism to provide sample-level concepts. For each , we define a binary vector by thresholding: . Since each corresponds to a latent trajectory , this acts as a coarse label for sample , indicating which of the underlying trajectories it is most influenced by. Next, we adapt our transformer to predicts both the binary pattern of and the final disease status, adding a concept-level loss that computes a logistic loss coordinatewise (see Listing 1). The results are displayed in Table 2. Notably, using this concept definition, the model performs better than the original, unconstrained transformer, suggesting closer alignment to the original generative mechanism. In general, better CBMs require denser concept annotation. Note, however, that recent work has proposed alternatives that discover concepts de novo or identify cheap concept proxies following model training (Yuksekgonul et al., 2023; Ghorbani et al., 2019).

### v2

@@ 2.2.1 Sparse logistic regression
where . This optimization identifies a sparse coefficient vector that maximizes the log-likelihood of pairs . The parameter controls sparsity: higher values yield simpler but less expressive models. The combination of sparsity and linearity ensures interpretability. Sparsity allows many features within to be ignored during prediction, while linearity ensures that when features do contribute, they enter the model transparently. Specifically, a one-unit increase in the -coordinate changes the relative risk for class by .
@@ 2.2.2 Decision trees
In contrast, decision trees do not solve a single optimization problem. Instead, they are defined through a recursive algorithm. Initially, we scan all dimensions of and determine thresholds such that maximizes classification accuracy. From these candidates, we select the optimal , defining the first partition along the axis . The two elements of the partition correspond to two leaves of a tree split according to . Next, for a tree with leaves, we iteratively split each element along axis at a threshold , for all pairs of leaves and dimensions , selecting the split that maximizes accuracy. This process stops once model complexity outweighs performance on a holdout set.
Interpretability here stems from parsimony and simulatability. The final tree has only as many partition elements as leaves, and with sufficient cost on model complexity, this can be easily visualized as a single tree. Further, the axis-aligned splits yield straightforward descriptions for each partition element, recoverable by tracing a path from the root to the leaf. Finally, predictions remain constant within each element, changing only across partition boundaries.
@@ 2.2.3 Advances
Sparse and tree-based models are the foundation for many proposals for interpretable machine learning (Xin et al., 2022; Zhong et al., 2023). For example, Zeng et al. (2016) introduced fast decision trees that match black box models in recidivism prediction, a domain with significant moral ramifications where interpretability is crucial. Similarly, Hazimeh et al. (2020) designed a decision tree variant that competes effectively with deep learning in various operations research tasks. By leveraging advances in combinatorial optimization, their approach allows interpretable trees to be learned from very large datasets without excessive compute demands.
Similarly, -regularization is central to many modern interpretability workflows, even in problems that initially seem distant from traditional statistical learning frameworks. For example, Fridovich-Keil et al. (2022) developed an interpretable approach to the view synthesis problem, where many two-dimensional views of an object are consolidated into a coherent three-dimensional reconstruction. Instead of the typical deep encoder-decoder architecture, they used -regularized regression on the appropriate Fourier representation. This approach achieved better accuracy than competing deep learning approaches with far less sensitivity to hyperparameter selection, and could be trained at a fraction of the compute cost.
@@ 2.2.4 Application to the simulation
We next apply these intrinsically interpretable models to the simulated trajectories from Section 2.1. We explore two approaches. First, we concatenate time points for each participant into a single -dimensional vector. This method can uncover temporally localized predictive signals. Second, we derive time series summaries for each taxon, reducing the number of predictors to for different summaries. In our study, we focus solely on linear trends and curvature. For each subject and taxon , we fit a linear model , with time as a predictor. The estimated summarizes the linear trend for that trajectory. Similarly, we use the sum-of-squares of the temporal second differences to summarize curvature: . Note that summarization derives potentially relevant features that are not simple linear combinations of the original inputs.
@@
The holdout accuracy for the sparse logistic regression and decision tree models under these two representations of and across a range of sample sizes is given in Table 2. Regularization and tree complexity parameters were chosen with four-fold cross validation. For these data, the sparse logistic regression model generally outperforms the decision tree. Moreover, manually derived summary statistics yield better performance compared to concatenating all inputs. This suggests a general takeaway: a simple model on the appropriate representation can achieve good performance, and performance is contingent on the representation. The widely invoked “interpretability–accuracy tradeoff” overlooks this nuance. Further, Table 2 includes the training time for the models, illustrating a broad range of runtimes among interpretable models. Simple doesn’t always mean fast or scalable.
@@
With respect to interpretability, when , the sparse logistic regression model on concatenated inputs with the optimal used 37 taxon-by-time point features, while the decision tree needed 13 splits. In contrast, with the summary statistic featurization, the optimal sparse logistic regression model used 50 features – the problem is lower dimensional, but the signal is more enriched, a fact suggested by the difference in prediction performance. Surprisingly, this enriched signal is not accessible to the decision tree, which produces a tree with only two splits. These conclusions appear to hold across . For example, even when , the sparse logistic regressions on raw and featurized inputs selected 60 and 68 features, respectively.
Data | Model | Samples () | In-Sample Accuracy | Out-of-Sample Accuracy | Training Time (s) |
Original | Sparse Logistic Reg. | 0.5 | 81.2% | 78.0% | 3.9 |
1 | 82.1% | 81.0% | 15.9 | ||
5 | 78.1% | 77.9% | 22.9 | ||
10 | 77.5% | 77.0% | 37.83 | ||
Decision Tree | 0.5 | 91.6% | 70.2% | 95.9 | |
1 | 91.3% | 76.5% | 497.9 | ||
5 | 91.4% | 75.7% | 3255.3 | ||
10 | 90.3% | 78.4% | 7676.9 | ||
Transformer | 0.5 | 100% | 86.4% | 91.2 | |
1 | 90.7% | 76.0% | 246.6 | ||
5 | 96.6% | 86.4% | 1167.6 | ||
10 | 87.6% | 87.7% | 1964.4 | ||
Concept Bottleneck | 0.5 | 98.4% | 84.8% | 110.9 | |
1 | 99.2% | 73.2% | 211.6 | ||
5 | 90.4% | 88.2% | 892.8 | ||
10 | 92.5% | 89.0% | 1854.0 | ||
Featurized | Sparse Logistic Reg. | 0.5 | 92.6% | 87.8% | 0.4 |
1 | 88.8% | 85.4% | 8.1 | ||
5 | 86.5% | 85.3% | 32.7 | ||
10 | 85.9% | 85.5% | 119.5 | ||
Decision Tree | 0.5 | 74.4% | 69.6% | 1.8 | |
1 | 87.9% | 70.6% | 14.7 | ||
5 | 90.2% | 75.1% | 185.5 | ||
10 | 88.1% | 77.7% | 770.8 |
@@
The parsimony of these estimates aligns with our discussion of the characteristics of intrinsically interpretable models. In both approaches, prediction relies only on a small subset of the original inputs. But how much can we trust these features? Fig 2e - f shows the estimated for nonzero features obtained from separately fitting the sparse logistic regression approaches on independent splits. When using concatenated inputs, only two features overlap, while with summary statistics, this increases to 18. Therefore, while the model on raw time points may be more parsimonious, it is less stable. In this sense, despite its parsimony, the first model does not lend itself naturally to interpretation. Conversely, the second appears more accurate and stable. Moreover, its features, like curvature and trends, match the existing vocabulary of domain experts, making it the more interpretable model in this context. The larger lesson is that interpretability refers to a constellation of model properties – parsimony and simulatability but also accuracy, stability, and contextual relevance – and it is worthwhile to think in these more refined terms rather than simply “interpretable” and “black box” models.
@@ 2.3 Deep learning and XAI
The sparse logistic regression with derived summary statistics seems to be our best option in the case study. However, we should be concerned by the fact that we manually define our features – what other potentially predictive features might we be missing out on? Black box models are appealing because they automate the representation learning process (Bengio, 2009; Bengio et al., 2013). For example, in computer vision, deep learning models can identify higher-order, semantically meaningful features from raw pixel inputs (Zeiler and Fergus, 2013; Simonyan et al., 2014; Yosinski et al., 2015; Achtibat et al., 2023; Bykov et al., 2023). The can learn features that activate on images of eyes simply by being given training data of faces and an appropriate loss function (Ross et al., 2017). Applying a deep learning model to our microbiome data could reveal predictive temporal and taxonomic features (e.g., the presence of particular subcommunities) that we had not previously considered. The challenge, then, would be to explain these models in a way that advances scientific knowledge and ensures accurate medical diagnoses. To demonstrate the process, let’s explore using a transformer encoder model in our simulation.
@@ 2.3.1 Transformers
Specifically, consider a sequence of tokens . These tokens can be embedded into a sequence of high-dimensional (e.g., 512) real-valued vectors. In a word-based language model, could be one-hot encodings of the words from the language’s vocabulary (Pennington et al., 2014), while in a vision problem, might represent neighboring patches from an image (Kolesnikov et al., 2021). The vectors are meant to capture the high-level features of the original token (word or patch) within its larger context (sentence or image). Therefore, they are much more valuable for downstream analysis than word identity or pixel values considered in isolation.
@@ 2.3.2 Embedding visualization
For example, Coenen et al. (2019) used embeddings to analyze the BERT language model (Devlin et al., 2019). In one experiment, they examined sentences containing words that can have multiple meanings depending on context, like, “fair.” By computing embeddings for instances of “fair” within these sentences, it was observed that although the word remained the same, its embeddings distinctly clustered in a two-dimensional principal components projection. This clustering reflected the context dependence of the word’s meaning, with clusters representing the legal (“fair use”), mathematical (“fair coin”), and civic (“world’s fair”) uses. This qualitative study can also be accompanied by quantitative analysis. For example, Coenen et al. (2019) and Hewitt and Manning (2019) used linear probes to quantify the extent to which grammatically structure (represented by parse tree representations) could be reconstructed from BERT embeddings. Figure 3 in Section 2.2.4 gives embedding visualization examples in the context of our simulation study. In particular, Figure 3d parallels the approach of Coenen et al. (2019), focusing on a species instead of a word.
@@ 2.3.3 Integrated gradients
When seeking local explanations, embedding visualization lacks sufficient specificity. We often have questions in mind that require more than model-based summaries – for example, what was it about a hotel review that led to its classification as negative sentiment, or how could a loan applicant modify their application so that it can be approved? Addressing these questions requires thinking carefully about the model’s sensitivity to perturbation. Many methods formalize this intuition – influence functions (Koh and Liang, 2017), Shapely values (Lundberg and Lee, 2017; Fumagalli et al., 2023), LIME (Ribeiro et al., 2016), GradCAM (Selvaraju et al., 2017), and Integrated Gradients (IG) (Sundararajan et al., 2017) for example. Here, we dive into the mechanics of IG.
@@
where is a reference sample (e.g., an all-black “empty” image) and is the model’s predicted probability for the input being assigned to class . The gradient represents how infinitesimal changes in each coordinate of affects the probability of assigning the input to class . At first, this appears to be sufficient for local explanation – large gradient coordinates correspond to words or pixels with the largest influence on the class assignment. However, this often fails in practice due to saturation. Specifically, deep learning classifiers often use a multiclass logistic function in the final layer, and for large input magnitudes, this function’s gradients approach zero, rendering explanations based on unsatisfactory.
@@
Integrated gradients and its local explanation counterparts are widely used in XAI, but they are also among the most controversial (Lundstrom et al., 2022). Both theoretical and empirical results suggest that explanations do not necessarily respect expectations. For example, Adebayo et al. (2018) observed that saliency maps, including those from IG, derived from randomly initialized neural network classifiers look closely resemble those from properly trained ones. A theoretical study by Bilodeau et al. (2024) found that locally similar functions could nonetheless be assigned different explanations at their query points, indicating that global properties can influence local explanations. These alarming findings have renewed interest in the interpretability community to develop local explainability techniques with formal guarantees.
@@ 2.3.4 Concept bottleneck models
The explainability techniques we have discussed so far depend only on the model and either the original training data, for embedding analysis, or the specific example to be explained, for integrated gradients. This is both a strength and a limitation. On the one hand, it ensures generality, the same explanation method can be effective across various applications. On the other, it limits the incorporation of domain-specific language into the explanation. Concept Bottleneck Models (CBMs) (Koh et al., 2020) offer an abstraction for infusing domain-specific language into models. CBM “concepts” bridge model and human representations of relevant data features. The overarching goal of this class of models is to create representations that can be manipulated by both the model (as high-dimensional, numerical vectors) and experts (as community-derived concepts).
@@
On a test sample , predictions are made using and . Note that ground truth concept labels are not needed at this stage. Typically, a simple model , like multiclass logistic regression, is used. CBMs automate human-interpretable feature extraction, bridging manual feature engineering with black box representation learning. While manually designing a “wing color” feature extractor would be tedious, it is helpful to know that the model learns a proxy for such a feature, and its output can be used like any human engineered feature. Moreover, this approach streamlines counterfactual reasoning. For example, we can ask what the model would have predicted had the wing been yellow instead of red, simply by intervening on the coordinate of associated with wing color. This is important in fairness or recourse analysis (Mehrabi et al., 2021; Karimi et al., 2022). For example, if we know that a feature should not be used for prediction (e.g., the race of a loan applicant), then we can zero out the coefficient from that concept in the model .
@@ 2.3.5 Application to the simulation
With these XAI techniques, we can apply black box models to our microbiome case study and still draw scientific insight from them. The simulated dataset is well-suited to a transformer encoder architecture, where each sample corresponds to a word and the full trajectory to a sentence. The final healthy vs. disease classification is analogous to text sentiment classification. Guided by these parallels, for each we trained a GPT-2 architecture with a random sample of simulated participants; the remaining are reserved for validation. We used default hyperparameters from a public implementation with the two modifications. First, we changed the input dimension to 144 to one-hot encode the species. Second, we reduced the number of layers to 6 from the default of 12, since the sample sizes are relatively small. After training this model for 70 epochs, the model achieves a holdout accuracy ranging from 76.0% () to 87.7% (). This performance is noteworthy considering that in several cases it exceeds the best models with handcrafted features (cf. Table 2), which were designed with the true generative mechanism in mind. Indeed, considering the close alignment between handcrafted features and the generative mechanism, we suspect that the Bayes error for this problem is between 10 and 15%. The ability to match or exceed a handcrafted model underscores the appeal of modern deep learning. It is not necessarily about achieving better performance than interpretable counterparts on all problems; rather, it comes from the potential to bypasses feature engineering – a traditionally labor-intensive and time-consuming step – while achieving comparable performance. Moreover, this simulation highlights the attractive scaling of deep learning models in both out-of-sample accuracy and running time, since performance continues to improve with larger sample sizes. This makes them valuable in applications where frequent errors cannot be tolerated, even if they could be explained. This is often the case in technological systems that have to operate well even with limited human oversight, like flight control (Chai et al., 2021) or online recommender systems (Resnick and Varian, 1997). Indeed, in these settings, reliability and robustness may be more important than interpretability alone.
@@
The remaining challenge lies in understanding how the models achieved their performances. We consider each of the techniques discussed above, focusing on the models trained with samples. Fig 3 offers two global views of the data. Fig 3a gives the sparse PCA projections derived from the scaled matrix of community profiles over time, while Fig 3b shows the analog for the transformer representations. The embeddings in (b) more clearly distinguish healthy and diseased groups, capturing more of the total variance within the top two dimensions. This is expected, since the final classification uses a linear hyperplane, encouraging the transformer to learn a linearly separable representation. To better understand the features encoded by the transformer, we focus on the subset of embeddings associated with species 21. Fig 3c shows sparse PCA applied to these embeddings. We have highlighted samples that lie near the linear interpolation between samples 110 and 378 in the original embedding space; their original trajectories are shown in panel (d). Along this interpolation, the peaks and valleys become less pronounced, suggesting that the model has learned to differentiate curves with different trajectory properties. This is a subtle feature, and one that we know by design is related to the response. In contrast, some visually prominent features unrelated to disease, like the maximum value of the series, are appropriately disregarded.
@@
We next study local explanations using IG. Based on the input structure, IG returns a value for each time point-by-taxon combination. Example values are given in Fig 4. In each panel, we can identify time points that informed the final classification and the direction in which perturbations change the target class label. Blue squares denote that an increased abundance at that time point increases the probability of correct classification; red circles suggests the opposite. Larger symbols imply larger changes in probability. For example, for subject 6, the first ten time points and the valley surrounding time 20 were most important. Increases in either of these intervals increase the probability of correctness. In contrast, for subject 63, an increase in the valley surrounding time 20 decreases this probability. Since these subjects are from opposite classes, deeper valleys around time 20 seem to be related to health. However, beware that these conclusions are essentially local, and this pattern need not hold across all samples for IG to declare it important. In general, IG effectively localizes the responsibility for a classified instance down to a subset of time points, aiding error analysis and uncovering subsets of features that could flip the class of a prediction.
@@
Next, we take a closer look at the use of CBMs in our case study. Recall that concepts are a form of dense annotation for each sample. We extend our original generative mechanism to provide sample-level concepts. For each , we define a binary vector by thresholding: . Since each corresponds to a latent trajectory, this acts as a coarse label for sample , indicating which of the underlying trajectories it is most influenced by. Next, we adapt our transformer to predict both the binary pattern of and the final disease status, adding a concept-level loss that computes a logistic loss coordinatewise (see Listing 1). The results are displayed in Table 2. Notably, using this concept definition, the model performs better than the original, unconstrained transformer, suggesting closer alignment to the original generative mechanism. In general, better CBMs require denser concept annotation. Note, however, that recent work has proposed variations that discover concepts de novo or identify cheap concept proxies following model training (Yuksekgonul et al., 2023; Ghorbani et al., 2019).